OpenStudy (anonymous):

find the indefinite integral of dx/8x^3

OpenStudy (anonymous):

$$\int\frac{dx}{8x^3}=\int\frac18x^{-3}\,dx=\frac18\int x^{-3}\,dx$$ Do you know the power rule?

OpenStudy (anonymous):

no not really i have my first lecture on it today and i didnt get to write down all the rules

OpenStudy (unklerhaukus):

if \(n,\alpha\) are both constants \[\int \alpha x^{n}\,\mathrm dx=\alpha \frac{x^{n+1}}{n+1}+c\]